Grade 11Algebra

15 men can complete work in 210 days. They started work but at end of 10 days, 15 additional men, with double efficiency, were inducted. How many days, did they all take to finish the work?

Girish Verma
10 Years agoGrade 11

4 Answers

G Maheshbabu
10 Years ago
the men took 110 days to complete their work,when additionally 15 men are added.the 30 members are complete their work in 110 days
Rajendra Patra
9 Years ago
NOTE:-AFTER CANCELLING OF 10 DAYS WORK,THEN THE CAPACITY BETWEEN THE 15 MEN AND 15 ADDITIONAL MEN OF 200 DAYS WORK IS TO BE COMPARED.
 
SO,15 MEN\rightarrow210 DAYS
 
Rajendra Patra
9 Years ago
NOTE:-AFTER CANCELLING OF 10 DAYS,THE CAPACITY OF WORK DONE IS CALCULATED OF 200 DAYS BETWEEN THE 15 MEN AND 15 ADDITIONAL MEN OF DOUBLE EFFICIENCY.
SO,         15MEN \rightarrow210 DAYS
                \Rightarrow1MEN \rightarrow210× 15 DAYS
AT THE END OF 10 DAYS 1 MEN HAVE  TO DO WORK FOR \rightarrow(210-10)×15 DAYS
THEN, 15 ADDITIONAL MEN OF DOUBLE EFFICIENCY MEANS THEY CAN DO THE WORK OF  15×2  MEN’S WORK.
SO, THE TOTAL MEN AT THE END OF 10 BECOMES \rightarrow15+15×2=45 MEN.
LET’S COMPARE THE CAPACITY BETWEEN 15 MEN AND 45 MEN;
                                SO,15MEN AT THE END OF 10 DAYS WILL DO WORK FOR 200 DAYS;
                                45 MEN WILL DO\rightarrow(200×15)/45 DAYS=66 ×2/3 DAYS
HENCE,TOTAL DAYS TAKEN TO COMPLETE THE WORK =(10+66×2/3) DAYS=76 2/3 DAYS ………………(ANS)
Manisha Nimsade
9 Years ago
15*200/45= 200/3 (210-10days =200 double the efficiency means 15*2 =30. Hence 15+30=45)So total days10+ 200/3=76 2/3